OpenStudy (anonymous):

Find the equation of the tangent line to the curve f(x)=ln(x^2) at the point x=e

OpenStudy (anonymous):

first off, note that \(\ln(x^2)=2\ln(x)\) the take the derivative

OpenStudy (anonymous):

what is the derivative of ln? e

OpenStudy (anonymous):

you don't want the derivative of \(ln(e)\) which is just 1, you want the derivative of \(2\ln(x)\)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

what is the derivative of \(\ln(x)\) ?

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

so the derivative of \(2\ln(x)\) is \(\frac{2}{x}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now to find the slope at \(x=e\) replace \(x\) by \(e\) and get a slope of \(\frac{2}{e}\)

OpenStudy (anonymous):

the point is \((x,2)\) and the slope is \(\frac{2}{e}\) find the equation of the line by using the point - slope formula

OpenStudy (anonymous):

y-2=2/e x

OpenStudy (anonymous):

where do u get the point (x,2)?